# Massive Higher-Spin Fields in the Fractional Quantum Hall Effect

### v1

@@ Abstract
Incompressibility plays a key role in the geometric description of fractional quantum Hall fluids. It is naturally related to quantum area-preserving diffeomorphisms and the underlying Girvin-MacDonald-Plazman algebra, which gives rise to an emergent non-relativistic massive spin-$2$ mode propagating in the bulk. The corresponding metric tensor can be identified with a nematic order parameter for the bulk states. In the linearised regime with a flat background, it has been shown that this mode can be described by a spin-$2$ Schroedinger action. However, quantum area-preserving diffeomorphisms also suggest the existence of higher-spin modes that cannot be described through nematic fractional quantum Hall states. Here, we consider p-atic Hall phases, in which the corresponding p-atic order parameters are related to higher-rank symmetric tensors. We then show that in this framework, non-relativistic massive chiral higher-spin fields naturally emerge and that their dynamics is described by higher-spin Schroedinger actions. We finally show that these effective actions can be derived from relativistic massive higher-spin theories in $2+1$ dimensions after taking a non-relativistic limit.
@@
Introduction: The notion of order parameter is useful as a unifying concept for describing states of matter with emergent macroscopic properties. For instance, superconductors [1] and liquid crystals [2] are characterised by local-order parameters, which reflect the presence of spontaneously broken symmetries. On the other hand, topological phases such as fractional quantum Hall (FQH) states are instead characterised by topological invariants. Consequently, they cannot be described within Landau’s theory of phase transitions [3, 4, 5]. However, there exist several states of matter where local order and topological phases can co-exist, such as topological superconductors [6, 7, 8, 9, 10, 11] and nematic FQH phases [12, 13, 14, 15, 16, 17, 18, 19, 20]. In the latter, nematicity can be induced, for instance, through a tilted magnetic field, which naturally stretches the Landau orbits while preserving the areas the orbits encircle. The corresponding nematic order, which is associated to a rank-$2$ symmetric tensor order parameter, plays a central role in the identification of an emergent “spin-$2$” collective massive mode that appears in the bulk [21, 22, 23, 24, 25, 26] and has been recently observed experimentally [27]. The importance of nematic order in this context can be qualitatively understood by noting that Hall nematic phases, which break rotational symmetry while preserving a $C_{2}$ discrete rotational symmetry, preserve the topological states while the corresponding deformations can be seen as special kinds of area-preserving deformations of the Hall fluids. The corresponding emergent quantum (non-commutative) geometry is, in fact, invariant under quantum area-preserving diffeomorphisms [28, 29, 30, 31], that encode the incompressibility of quantum Hall fluids, and consequently are responsible for the existence of an energy bulk gap. These features are encoded in the so called Girvin-MacDonald-Plazman (GMP) algebra [32], also known as $W_{\infty}$ algebra in the high-energy-physics literature [28] [^1] The $W_{\infty}$ algebra mentioned here is a deformation of the algebra of area-preserving diffeomorphisms of a cylinder, called the $w_{\infty}$ algebra. This deformation is linear in the generators and was constructed in [78]. This infinite-dimensional algebra should be distinguished from (i) the 3D higher-spin algebra $hs(1,1)$ that was used in [34] to construct a gauge theory of 3D massless higher spins and is related to the algebra of area-preserving diffeomorphims of the two-dimensional hyperbolic space and (ii) from the asymptotic symmetry algebra of the $hs(1,1)$ gauge theory which is a nonlinearly realised $W_{\infty}$ algebra [79] (see also [80, 81]). For a one-parameter supersymmetric extension of the (linearly realised) $W_{\infty}$ algebra, see [82].. Thus, the emergent geometry in Hall fluids is associated to a non-relativistic massive graviton-like mode, which at linearised level around a flat background can be described by a spin-$2$ Schroedinger action [25]. A characteristic feature of the GMP/$W_{\infty}$ algebra is that, besides a generator of conformal spin two, it also contains an infinite number of generators of increasing conformal spin, with each spin occurring once [34]. This suggests the existence of additional bosonic bulk modes of spin greater than two (aka higher-spin modes), and this expectation is supported by several theoretical and numerical evidences [35, 36, 37, 38]. However, the quantum-field-theory characterization of these massive modes remains unclear.
@@
In this work, we partially address this important open question by employing a complementary approach in which we show that these higher-spin modes can naturally appear in the fractional quantum Hall effect (FQHE) via p-atic phases, with p even such that inversion symmetry is preserved [39]. In p-atic liquid crystals, the corresponding p-atic order parameters are higher-rank traceless symmetric tensors, which in two spatial dimensions are related to $C_{\textrm{p}}$ symmetries [40, 41]. They are well known in soft matter and naturally generalize nematic phases [2, 42, 43, 44, 45, 46]. Notice that although several p-atic Hall phases have been already proposed and studied [47, 48, 49, 50, 51], their relation with higher-spin fields has been overlooked. Here, we will show that within this framework, massive higher-spin fields naturally emerge and their dynamics is described by higher-spin Schroedinger actions near the isotropic phase transition. In particular, we will focus on the chiral and massive spin-$4$ mode, which has been numerically discussed in Ref. [37] and that we will show to be related to a tetratic order parameter. We will finally show, through symmetry and field-theory arguments, that these effective actions can be derived from relativistic massive higher-spin theories after taking a suitable non-relativistic limit.
@@
P-atic phases in the FQHE: In this section, firstly we summarize the effective-field-theory description of nematic order in the FQHE by mainly following Ref. [14]. Secondly we generalize those results to the case of tetratic Hall states. De Gennes’ nematic order in liquid crystals is characterised by a rank-$2$ symmetric and traceless order parameter, which in two spatial dimensions takes the form $Q_{ab}({\bf{x}},t)=Q_{N}(2d_{a}d_{b}-\delta_{ab})$, where ${\bf{d}}=(\cos\vartheta/2,\sin\vartheta/2)$ is the planar nematic director [52]. Since it has only two independent components, we can define from it the complex order parameter
@@
with $A_{\mu}$ and $a_{\mu}$ the electromagnetic and U(1) emergent gauge fields, respectively, and where $\mu,\nu\in\{0,1,2\}$. Here, $q$ is an odd-integer number associated to the filling factor such that the Hall conductivity is quantised to $(1/q)(e^{2}/h)$. The corresponding effective action for the nematic Laughlin states is given by [14]
@@
where $\dot{Q}_{ab}\equiv\partial Q_{ab}/\partial t$, $a^{N}_{0}$ can be associated to a Berry phase, while $C$ is a constant parameter (it is related to the equal-time structure factor as shown in Ref. [14]). Close to the phase transition from an isotropic to a nematic FQH phase, $\mathcal{L}_{N}$ in the above equation identifies the effective Landau-de Gennes (LdG) Lagrangian
@@
We are now ready to discuss the generalization of the above results to the case of tetratic order. The latter refers to a generalizion of nematic order, in which a fourfold rotational symmetry $C_{4}$ replaces the $C_{2}$ of nematic phases. This phase can be either considered as the result of a deformation of the magnetic field or of the confining potential preserving the $C_{4}$ symmetry, or as a result of local orientational ordering introduced by interactions, possibly supplemented by additional p-atic modes [47, 48, 49, 50, 51]. Being inversion invariant, tetratic phases are compatible with FQH states [39]. The tetratic order parameter is a symmetric traceless rank-$4$ tensor $Q_{abcd}$. Importantly, in two space dimensions, like for the nematic order parameter, the traceless symmetric tensor $Q_{abcd}=Q_{T}(8d_{a}d_{b}d_{c}d_{d}-8\delta_{(ab}d_{c}d_{d)}+\delta_{(ab}\delta_{cd)})$, where ${\bf{d}}=(\cos\vartheta/4,\sin\vartheta/4)$ is the planar tetratic director, admits only two independent components that can be used to define the complex order parameter
@@
This peculiar feature that holds only in two dimensions allows us to generalize most of the results derived in Ref. [14] in the long-wavelength limit. In fact, also in the current case, we can define an action for the tetratic Laughlin states, which reads
@@
where $G$ is a constant parameter. Close to the transition from an isotropic to a tetradic FQH phase, the corresponding effective LdG Lagrangian $\mathcal{L}_{T}$ [40, 41] augmented by the Berry phase term $a^{T}_{0}$ is given by
$\mathcal{L}_{T}=\frac{k^{\prime}_{1}}{8}\epsilon_{bd}Q_{afgb}\dot{Q}_{afgd}-\frac{k^{\prime}_{2}}{8}(\partial_{d}Q_{abcd})^{2}-\frac{k^{\prime}_{3}}{2}(Q_{abcd})^{2},$ (8)
@@
and the corresponding spectrum identifies a gapped spin-$4$ collective mode. All the results presented above can be naturally generalised to spin-p modes associated to $C_{\textrm{p}}$ symmetries and p-atic Hall states. See also Ref. [36] for a complementary way to introduce effective bulk higher-spin fields in the FQHE, that however does not seem to be directly linked to p-atic phases.
From the Fierz-Pauli equations to higher-spin Schroedinger actions using symmetries: Before discussing the spin-$4$ and higher-spin massive modes, we first show how the LdG effective-field-theory description of the GMP mode can be obtained by concentrating on symmetries only [53]. Our starting point are the equations of motion for a free relativistic massive spin-$2$ particle of mass $m$
@@
with $S_{ab}$ symmetric and traceless. We then observe that under a Lorentz boost a time index transforms to a spatial index but that the inverse transformation from a spatial index to a time index is suppressed by a factor of $1/c^{2}$. This implies that, after taking the limit $c\to\infty$, the boost becomes a transformation that only acts in one direction on the different components of $h_{\mu\nu}$ as follows
@@
From this we deduce that, unlike the $\phi$ and $v_{a}$ components, the symmetric traceless field $S_{ab}$ forms by itself a representation of the contracted Poincaré group which is the Galilei group. In $2+1$ dimensions one can combine the two independent components $S_{11}$ and $S_{12}$ of $S_{ab}$ into the complex field $\mathcal{S}_{N}=S_{11}+iS_{12}$, that transforms under infinitesimal spatial rotations generated by $\Lambda_{ab}=\epsilon_{ab}\lambda$ not as a scalar but as
@@
The second term on the r.h.s. represents an emerging U(1) symmetry. Furthermore, we find that $\mathcal{S}(x)$ transforms under internal parity (P) $x^{1}\to-x^{1}$, time-reversal (T) $t\to-t$ and inversion (I) $x^{1,2}\to-x^{1,2}$ as follows
@@
$\delta\mathcal{Q}_{N}=-(\Lambda^{a}t)\partial_{a}\mathcal{Q}_{N}+im\Lambda_{a}x^{a}\mathcal{Q}_{N}.$ (16)
Furthermore, in the limit $c\to\infty$ the symmetry under parity P and time-reversal T is broken. This is due to the fact that both the $t\to-t$ time-reversal transformation and the complex conjugation that occurs in the parity transformation (14) have the effect that when taking the limit of the transformation rule of the non-relativistic field $\mathcal{Q}_{N}$, the phase factor does not drop out which gives an invalid limit. Instead, we find that the inversion symmetry I is preserved. There exists a unique action invariant under translations, spatial rotations together with the modified boost of Eq. (16), that is of first order in the time derivatives, thus breaking the invariance under time-reversal symmetry. It is given by the spin-$2$ planar Schroedinger action
@@
This action corresponds to the complex rewriting of Eq. (4), thus allowing one to interpret $\mathcal{Q}_{N}$ as the nematic order parameter (1). Whereas the FP equations (10) describes two helicity modes, that under parity are mapped to each other, the Schroedinger action (17) describes just a single mode and parity is broken. The reason that we lost a degree of freedom is that instead of complexifying all FP fields, we combined two real FP fields into a single complex field. In summary, we conclude that the LdG description of the fluctuations of the nematic phase order parameter around the isotropic phase in the long wavelength limit given in Eq. (4) can be recovered from a non-relativistic limit of the three-dimensional FP theory.
We can now generalize the above discussion to the spin-$4$ case, which is the next higher-spin field that preserves inversion symmetry. We start from a rank-$4$ symmetric tensor field $\Phi_{\mu\nu\rho\sigma}(x)$ that satisfies

### v2

@@ Abstract
Incompressibility plays a key role in the geometric description of fractional quantum Hall fluids. It is naturally related to quantum area-preserving diffeomorphisms and the underlying Girvin-MacDonald-Plazman algebra, which gives rise to an emergent non-relativistic massive spin-$2$ mode propagating in the bulk. The corresponding metric tensor can be identified with a nematic order parameter for the bulk states. In the linearized regime with a flat background, it has been shown that this mode can be described by a spin-$2$ Schroedinger action. However, quantum area-preserving diffeomorphisms also suggest the existence of higher-spin modes that cannot be described through nematic fractional quantum Hall states. Here, we consider p-atic Hall phases, in which the corresponding p-atic order parameters are related to higher-rank symmetric tensors. We then show that in this framework, non-relativistic massive chiral higher-spin fields naturally emerge and that their dynamics is described by higher-spin Schroedinger actions. We finally show that these effective actions can be derived from relativistic massive higher-spin theories in $2+1$ dimensions after taking a non-relativistic limit.
I Introduction
@@
The notion of order parameter is useful as a unifying concept for describing states of matter with emergent macroscopic properties. For instance, superconductors [1] and liquid crystals [2] are characterized by local-order parameters, which reflect the presence of spontaneously broken symmetries. On the other hand, topological phases such as fractional quantum Hall (FQH) states are instead characterized by topological invariants. Consequently, they cannot be described within Landau’s theory of phase transitions [3, 4, 5]. However, there exist several states of matter where local order and topological phases can co-exist, such as topological superconductors [6, 7, 8, 9, 10, 11] and nematic FQH phases [12, 13, 14, 15, 16, 17, 18, 19, 20]. In the latter, nematicity can be induced, for instance, through a tilted magnetic field, which naturally stretches the Landau orbits while preserving the areas the orbits encircle. The corresponding nematic order is associated to a rank-$2$ symmetric tensor order parameter that gives an emergent “spin-$2$” collective massive mode that appears in the bulk [21, 22, 23, 24, 25, 26] and has been recently observed experimentally [27]. The importance of nematic order in this context can be qualitatively understood by noting that Hall nematic phases, which break rotational symmetry while preserving a $C_{2}$ discrete rotational symmetry, preserve the topological states while the corresponding deformations can be seen as special kinds of area-preserving deformations of the Hall fluids. The corresponding emergent quantum (non-commutative) geometry is, in fact, invariant under quantum area-preserving diffeomorphisms [28, 29, 30, 31], that encode the incompressibility of quantum Hall fluids, and consequently are responsible for the existence of an energy bulk gap. These features are encoded in the so called Girvin-MacDonald-Plazman (GMP) algebra [32], also known as $W_{\infty}$ algebra in the high-energy-physics literature [28] [^1] The $W_{\infty}$ algebra mentioned here is a deformation of the algebra of area-preserving diffeomorphisms of a cylinder, called the $w_{\infty}$ algebra. This deformation is linear in the generators and was constructed in [86]. This infinite-dimensional algebra should be distinguished from (i) the 3D higher-spin algebra $hs(1,1)$ that was used in [34] to construct a gauge theory of 3D massless higher spins and is related to the algebra of area-preserving diffeomorphisms of the two-dimensional hyperbolic space and (ii) from the asymptotic symmetry algebra of the $hs(1,1)$ gauge theory which is a nonlinearly realized $W_{\infty}$ algebra [87] (see also [88, 89]). For a one-parameter supersymmetric extension of the (linearly realized) $W_{\infty}$ algebra, see [90].. Thus, the emergent geometry in Hall fluids is associated to a non-relativistic massive graviton-like mode, which at linearised level around a flat background can be described by a spin-$2$ Schroedinger action [25]. A characteristic feature of the GMP/$W_{\infty}$ algebra is that, besides a generator of conformal spin two, it also contains an infinite number of generators of increasing conformal spin, with each spin occurring once [34]. This suggests the existence of additional bosonic bulk modes of spin greater than two (aka higher-spin modes), and this expectation is supported by several theoretical and numerical evidences [35, 36, 37, 38]. However, the quantum-field-theory characterization of these massive modes remains unclear.
@@
In this work, we partially address this important open question by employing a complementary approach in which we show that these higher-spin modes can naturally appear in the fractional quantum Hall effect (FQHE) via p-atic phases, with p even such that inversion symmetry is preserved [39]. In p-atic liquid crystals, the corresponding p-atic order parameters are higher-rank traceless symmetric tensors, which in two spatial dimensions are related to $C_{\textrm{p}}$ symmetries, namely p-fold rotational symmetries [40, 41]. They are well known in soft matter and naturally generalize nematic phases [2, 42, 43, 44, 45, 46]. Notice that although several p-atic Hall phases have been already proposed and studied [47, 48, 49, 50, 51], their relation with higher-spin fields has been overlooked. Here, we will show that within this framework, massive higher-spin fields naturally emerge and their dynamics is described by higher-spin Schroedinger actions near the isotropic phase transition. In particular, we will focus on the chiral and massive spin-$4$ mode, which has been numerically discussed in [37] and that we will show to be related to a tetratic order parameter. We will also show, through symmetry and field-theory arguments, that these effective actions can be derived from relativistic massive higher-spin theories after taking a suitable non-relativistic limit. Finally, we will propose a measurement protocol to detect spin-4 modes in FQH states.
II P-atic phases in the FQHE
@@
In this section, firstly we summarize the effective-field-theory description of nematic order in the FQHE by mainly following [14]. Secondly, we generalize those results to the case of tetratic Hall states. De Gennes’ nematic order in liquid crystals is characterized by a rank-$2$ symmetric and traceless order parameter, which in two spatial dimensions takes the form $Q_{ab}({\bf{x}},t)=Q_{N}(2d_{a}d_{b}-\delta_{ab})$, where ${\bf{d}}=(\cos\vartheta/2,\sin\vartheta/2)$ is the planar nematic director [52]. Since it has only two independent components, we can define from it the complex order parameter
@@
with $A_{\mu}$ and $a_{\mu}$ the electromagnetic and U(1) emergent gauge fields, respectively, and where $\mu,\nu\in\{0,1,2\}$. Here, $q$ is an odd integer number associated with the filling factor such that the Hall conductivity is quantized to $(1/q)(e^{2}/h)$. The corresponding effective action for the nematic Laughlin states is given by [14]
@@
where $\dot{Q}_{ab}\equiv\partial Q_{ab}/\partial t$, $a^{N}_{0}$ can be associated with a Berry phase, while $C$ is a constant parameter (it is related to the equal-time structure factor as shown in [14]). Close to the phase transition from an isotropic to a nematic FQH phase, $\mathcal{L}_{N}$ in the above equation identifies the effective Landau-de Gennes (LdG) Lagrangian
@@
We are now ready to discuss the generalization of the above results to the case of tetratic order. The latter refers to a generalization of nematic order, in which a fourfold rotational symmetry $C_{4}$ replaces the $C_{2}$ of nematic phases. This phase can be either considered as the result of a deformation of the magnetic field or of the confining potential preserving the $C_{4}$ symmetry, or as a result of local orientational ordering introduced by interactions, possibly supplemented by additional p-atic modes [47, 48, 49, 50, 51]. Being inversion invariant, tetratic phases are compatible with FQH states [39]. The tetratic order parameter is a symmetric traceless rank-$4$ tensor $Q_{abcd}$. Importantly, in two space dimensions, like for the nematic order parameter, the traceless symmetric tensor $Q_{abcd}=Q_{T}(8d_{a}d_{b}d_{c}d_{d}-8\delta_{(ab}d_{c}d_{d)}+\delta_{(ab}\delta_{cd)})$, where ${\bf{d}}=(\cos\vartheta/4,\sin\vartheta/4)$ is the planar tetratic director, admits only two independent components that can be used to define the complex order parameter
@@
This peculiar feature that holds only in two dimensions allows us to generalize most of the results derived in [14] in the long-wavelength limit. In fact, also in the current case, we can define an action for the tetratic Laughlin states, which reads
@@
where $G$ is a constant parameter. Close to the transition from an isotropic to a tetratic FQH phase, the corresponding effective LdG Lagrangian $\mathcal{L}_{T}$ [40, 41] augmented by the Berry phase term $a^{T}_{0}$ is given by
$\mathcal{L}_{T}=\frac{k^{\prime}_{1}}{8}\epsilon_{bd}Q_{afgb}\dot{Q}_{afgd}-\frac{k^{\prime}_{2}}{8}(\partial_{f}Q_{abcd})^{2}-\frac{k^{\prime}_{3}}{2}(Q_{abcd})^{2},$ (8)
@@
that can be diagonalized in momentum space
$\mathcal{H}_{T}=\int\frac{d^{2}k}{(2\pi)^{2}}E_{\bf{k}}\gamma_{\bf{k}}^{\dagger}\gamma_{\bf{k}},\hskip 1.70709pt[\gamma_{\bf{k}},\gamma^{\dagger}_{\bf{k^{\prime}}}]=(2\pi)^{2}\delta({\bf{k-k^{\prime}}}),$ (10)
where $\gamma_{\bf{k}}^{\dagger}$ are bosonic creation operators and $E_{\bf{k}}=m^{\prime}+({k^{\prime}_{2}}/4){\bf{k}}^{2}$, where $m^{\prime}=k^{\prime}_{3}$ is a constant mass term. Thus, the mode spectrum identifies a gapped spin-$4$ collective mode. All the results presented above can be naturally generalized to spin-p modes associated to $C_{\textrm{p}}$ symmetries and p-atic Hall states. See also [36] for a complementary way to introduce effective bulk higher-spin fields in the FQHE, that is however motivated by area-preserving diffeomorphisms rather than by p-atic phases.
III From Fierz-Pauli equations to higher-spin Schroedinger actions using symmetries
The LdG effective-field-theory description of the GMP mode in (4) was first proposed in [53]. We now re-derive it by concentrating on symmetries only, before using the same approach to derive the spin-4 LdG effective theory (8) and discuss how to encompass also massive modes of higher spin.
A Nematic order parameter
Our starting point are the equations of motion for a free relativistic spin-$2$ particle of mass $m$:
@@ A Nematic order parameter
with $S_{ab}$ symmetric and traceless. We then observe that under a Lorentz boost a time index transforms to a spatial index but that the inverse transformation from a spatial index to a time index is suppressed by a factor of $1/c^{2}$ (see Eq. (34)). This implies that, after taking the limit $c\to\infty$, the boost becomes a transformation that only acts on the different components of $h_{\mu\nu}$ as follows
@@
From this we deduce that, unlike the $\phi$ and $v_{a}$ components, the symmetric traceless field $S_{ab}$ forms by itself a representation of the Galilei group, which is obtained as a $c\to\infty$ contraction of the Poincaré group. In $2+1$ dimensions one can combine the two independent components $S_{11}$ and $S_{12}$ of $S_{ab}$ into the complex field $\mathcal{S}_{N}=S_{11}+iS_{12}$, that transforms under infinitesimal spatial rotations generated by $\Lambda_{ab}=\epsilon_{ab}\lambda$ not as a scalar but as
@@
Furthermore, we find that $\mathcal{S}(x)$ transforms under internal parity (P) $x^{1}\to-x^{1}$, time-reversal (T) $t\to-t$ and inversion (I) $x^{1,2}\to-x^{1,2}$ as follows
@@
$\delta\mathcal{Q}_{N}=-(\Lambda^{a}t)\partial_{a}\mathcal{Q}_{N}+im\Lambda_{a}x^{a}\mathcal{Q}_{N},$ (17)
where the last new term leads to a Bargmann U(1) central extension. Furthermore, in the limit $c\to\infty$ the symmetry under parity P and time-reversal T is broken. This is due to the fact that both the $t\to-t$ time-reversal transformation and the complex conjugation that occurs in the parity transformation (15) leaves a factor $e^{-imc^{2}}$ in the transformation rule of the non-relativistic field $\mathcal{Q}_{N}$ which does not have a well defined $c\to\infty$ limit. Instead, we find that the inversion symmetry I is preserved. There exists a unique action invariant under translations, spatial rotations together with the modified boost of Eq. (17), that is of first order in the time derivatives, thus breaking the invariance under time-reversal symmetry. It is given by the spin-$2$ planar Schroedinger action
@@
This action corresponds to the complex rewriting of Eq. (4), thus allowing one to interpret $\mathcal{Q}_{N}$ as the nematic order parameter (1). Whereas the FP equations (11) describe two helicity modes, that under parity are mapped to each other, the Schroedinger action (18) describes just a single mode and parity is broken. In the following section we will also derive the Schroedinger action (18) from a direct $c\to\infty$ limit of a relativistic action, thus showing how the second helicity mode drops out in the non-relativistic limit, consistently with the previous discussion. In summary, we conclude that the LdG description of the fluctuations of the nematic phase order parameter around the isotropic phase in the long wavelength limit given in Eq. (4) can be recovered from a non-relativistic limit of the three-dimensional FP theory.
B Tetratic order parameter
We can now generalize the above discussion to the spin-$4$ case, which is the next higher-spin field that preserves inversion symmetry. We start from a rank-$4$ symmetric tensor field $\Phi_{\mu\nu\rho\sigma}(x)$ that satisfies the spin-4 FP equations [55]